# ERA-IT: Aligning Semantic Models with Revealed Economic Preference for Real-Time and Explainable Patent Valuation

### v1

@@ Abstract
Valuing intangible assets under uncertainty remains a critical challenge in the strategic management of technological innovation due to the information asymmetry inherent in high-dimensional technical specifications. Traditional bibliometric indicators, such as citation counts, fail to address this friction in a timely manner due to the systemic latency inherent in data accumulation. To bridge this gap, this study proposes the Economic Reasoning Alignment via Instruction Tuning (ERA-IT) framework. We theoretically conceptualize patent renewal history as a revealed economic preference and leverage it as an objective supervisory signal to align the generative reasoning of Large Language Models (LLMs) with market realities, a process we term Eco-Semantic Alignment. Using a randomly sampled dataset of 10,000 European Patent Office patents across diverse technological domains, we trained the model not only to predict value tiers but also to reverse-engineer the Economic Chain-of-Thought from unstructured text. Empirical results demonstrate that ERA-IT significantly outperforms both conventional econometric models and zero-shot LLMs in predictive accuracy. More importantly, by generating explicit, logically grounded rationales for valuation, the framework serves as a transparent cognitive scaffold for decision-makers, reducing the opacity of black-box AI in high-stakes intellectual property management.
@@ 1 Introduction
In the contemporary knowledge-driven economy, patents, as core intangible assets, are increasingly pivotal to firms’ strategic decisions and value creation ernst2003 . They serve as the fundamental currency of innovation, delineating a firm’s technological trajectory and securing market dominance hall2005 ; reitzig2004 . Accurate and timely patent valuation forms the bedrock of firms’ strategy for managing technological innovation  fischer2014 ; kalip2022 . However, achieving objective and accurate ex-ante valuation of patents remains a classic and persistent grand challenge in information science, information economics, and technological innovation management reitzig2004 . Figure 1 outlines the evolutionary path of patent valuation methodologies.
@@
The crux of this challenge lies in two facts. First, patents themselves are complex and information-dense hybrid "technical-legal" artifacts. Their true value is deeply encoded in unstructured technical texts han2015 . This creates profound information asymmetry between informed patent holders and external evaluators—such as investors and managers—who struggle to penetrate this “technical black box” and to transact efficiently in markets for technology akerlof1970 ; gambardella2007 ; gans2008 . This opacity creates severe market friction, where stakeholders struggle to distinguish high-value inventions from the vast noise of global filings, due to the high-dimensional complexity of mapping technical specifications to market outcomes pakes1986 ; trajtenberg1990 . Second, information asymmetry is inherent in the global patent ecosystem. Unlike domain-specific studies that benefit from homogeneous terminologies, a generalized valuation framework must contend with heterogeneous linguistic patterns across disparate technological sectors pakes1984 . In this context, traditional bibliometric indicators, such as citation counts, fail to serve as effective signals for timely decision-making due to systemic temporal latency; a patent typically accumulates meaningful citation data only years after its publication harhoff2003 . This "backward-looking" nature of bibliometrics prevents managers from assessing patent value in a way that is both real-time and closely reflects genuine market preferences. Consequently, there exists an urgent imperative to bridge the gap between technical disclosure and economic realization. This imperative drives the quest for advanced architectures capable of leveraging "forward-looking" signals, constituting a critical research frontier.
@@
In this context, historical data on patent renewal has gained theoretical prominence as a direct signal of "revealed economic preferences", since the act of renewal represents a "vote with the wallet" that is more closely aligned with market reality than any external proxy  pakes1986 ; pakes1989 ; schankerman1986 . From the perspective of economic signaling theory, the decision to pay monotonically increasing renewal fees reveals the patent holder’s preference, indicating that the asset’s private value exceeds its maintenance expenses kabore2012 ; kumar2024 . Unlike stochastic citations, renewal records provide a definitive, discrete signal of value maintenance available throughout the patent’s lifecycle. However, a fundamental methodological gap persists: how to effectively align this structured, low-dimensional economic-behavior signal with the high-dimensional, unstructured semantic information in patent texts to build an accurate and explainable predictive model. Existing frameworks have struggled to integrate this objective economic signal with the rich, unstructured semantic information embedded within patent documents arts2018 . Traditional Natural Language Processing (NLP) techniques often fail to capture the nuanced causal relationships between the specific exclusionary scope defined in the claims and the patent’s potential for economic longevity lee2020 .
@@
The recent advent of Large Language Models (LLMs) offers a transformative opportunity to overcome these linguistic barriers. LLMs possess the unique capacity for complex reasoning and for extracting latent semantic features. Nevertheless, applying general-purpose LLMs to patent valuation is non-trivial; these models often lack the domain-specific intuition required to align technical novelty with economic viability. To overcome this critical barrier in high-stakes Intellectual Property (IP) management in such contexts, there is an urgent need for adaptive technological and methodological innovations. The development of a predictive architecture that effectively fuses the objective ground truth of renewal data with the semantic reasoning capabilities of LLMs, a process that we term "Eco-Semantic Alignment”, represents a largely unexplored frontier in intelligent information systems zhang2022 .
@@
To bridge this methodological gap, this study proposes the Economic-Reasoning Alignment via Instruction Tuning (ERA-IT) framework, as illustrated in Figure 2. Our approach uses European Patent Office (EPO) renewal data as a supervisory signal to guide LLMs’ generative reasoning. Guided by a rigorous data fusion strategy, we operationalize the longitudinal payment of renewal fees as a value proxy. This allows us to translate economic behavior into a labeled dataset with discrete value tiers. This transforms the ambiguous task of valuation into a robust supervised learning problem. Subsequently, we employ a strategy that synergizes parameter-efficient fine-tuning with chain-of-thought prompting. This alignment enables the model to learn latent semantic patterns correlated with economic survival, effectively converting unstructured text into quantifiable predictive indicators.
Specifically, we explore three critical research questions underexplored in prior studies:
  • RQ1 (Efficacy): Does ERA-IT outperform state-of-the-art discriminative baselines and zero-shot LLMs in predicting patent value?
  • RQ3 (Generalizability): Is the framework robust across heterogeneous technological domains (International Patent Classification (IPC) sections) and low-resource settings?
To rigorously validate the efficacy and generalizability of this framework, we conducted a comprehensive empirical analysis using a dataset comprising over 10,000 EPO patents, randomly sampled across various IPC sections. This randomized sampling strategy ensures that our findings are not biased toward specific high-velocity domains and instead represent a robust solution applicable to the broader technological landscape. Our results demonstrate that this hybrid approach yields significantly superior predictive accuracy compared to conventional rule-based heuristics and statistical baselines arts2018 . By minimizing predictive latency in citation metrics, our system provides a scalable solution for automated high-precision IP valuation.
Our study pioneers a theoretical framework by aligning LLMs’ semantic reasoning with revealed economic preferences to establish a new paradigm for real-time and interpretable patent valuation. It has at least threefold research contributions, summarized as follows.
@@
  • Methodological Contribution in Causal Reasoning: We establish the algorithmic efficacy of the ERA-IT architecture, which synergizes parameter-efficient fine-tuning with domain-specific instruction tuning strategies. We provide empirical evidence that standard pre-trained models are insufficient for the specialized task of IP valuation, demonstrating that our domain-adaptation process captures the "Economic Chain-of-Thought" within obfuscated legal-technical texts.
@@ 2.1 Patent Valuation
Research on patent valuation has shifted from simple patent counts toward richer indicators that proxy the economic and innovative impact of inventions. Using the American Inventors Protection Act, which mandated earlier publication of most US patent applications, Hegde et al. (2023) show that earlier disclosure is associated with more and faster follow-on citations, fewer abandonments, and higher R&D investment, highlighting how publication rules shape both the private and social value of patents. Svensson (2022) analyzes four widely used patent value indicators, renewal length, forward citations, family size, and documents that each capture distinct facets of technological and economic value rather than a single underlying measure. Building on this indicator-based tradition, Ponta et al. (2021) propose the Innovation Patent Index (IPI), a composite innovation-performance measure that aggregates multiple patent characteristics related to technological quality, market potential, and scope into a firm-level index. Complementing these firm- and patent-level perspectives, Park et al. (2023) apply the previously introduced CD index of disruption to large-scale datasets of papers and patents, extend it with normalized variants and alternative measures, and show that science and technology have become less disruptive over time, illustrating the potential of advanced citation-based metrics for characterizing the qualitative role of inventions in the knowledge network.
However, from the perspective of ex-ante patent valuation, these approaches remain largely ex-post and rely on outcome variables that materialize only after a substantial time lag. Renewal behavior, composite indices such as IPI, and disruption-based citation metrics are typically calibrated on long-term renewal patterns, realized commercialization, or accumulated citation structures, and they make limited use of the full semantic content of patent texts at the level of individual inventions Svensson (2022); Ponta et al. (2021); Park et al. (2023). Moreover, these indicators generally summarize value at the firm or field level and do not provide explicit, stepwise economic rationales for why a particular patent should be assessed as more or less valuable at the time of filing. Because of these limitations, existing patent-valuation approaches remain constrained in their ability to provide economically grounded and explainable ex-ante value assessments for newly filed patents. These limitations underscore the need for ex-ante patent valuation frameworks that integrate semantic text analysis with interpretable reasoning, such as ERA-IT.
@@ 2.2 Semantic Analysis for IP Management
Recent work in IP management increasingly relies on semantic analysis of patent and related technical text. Ali et al. (2024) review the evolution of patent retrieval and group methods into families that range from keyword and vector space models to semantic and deep learning-based techniques for prior art search. Xu and Ono (2025) extend this line of work to industry–academia collaboration by proposing a semantic retrieval approach that uses text embeddings to analyze and rank collaborative research outputs. At the model level, Chikkamath et al. (2022) construct large-scale USPTO subclass datasets and investigate how a BERT for Patents encoder and strong non-neural baselines use claims and abstracts for multi-label patent classification. Complementary research on structured representations decomposes Chinese patent claims into function, structure, and purpose segments and applies pre-trained transformer-based models to classify them into technical categories, thereby improving classification accuracy and supporting retrieval of patents in similar technical domains Li et al. (2025). In parallel, Zini and Awad (2022) survey explainability techniques for deep NLP models, cataloguing model-agnostic and model-specific methods that interpret embeddings, internal representations, and output decisions for text-based tasks.
@@
Despite these advances, existing semantic approaches are optimised for retrieval, classification, or document alignment rather than for transparent ex-ante economic valuation of patents. The patent retrieval survey focuses on relevance and recall, leaving the relationship between semantic similarity and economic value largely implicit Ali et al. (2024). Semantic matching systems for industry–academia collaboration operate in high-dimensional embedding spaces but do not attempt to infer monetary or strategic value or to incorporate ground-truth financial signals Xu and Ono (2025). Models such as the BERT for Patents and the structured claim classifiers in the Chinese patent study are trained to predict technology codes, not renewal-based value tiers, and typically output numerical labels without explicit economic rationales Chikkamath et al. (2022); Li et al. (2025). Even where explainability is the primary topic, current surveys emphasise post-hoc interpretation tools for generic NLP tasks and note that rigorous evaluation in high-stakes legal and economic contexts remains limited Zini and Awad (2022). In contrast, ERA-IT explicitly fuses semantic reasoning with revealed preference renewal data and uses instruction-tuned chain-of-thought generation to map claim language to value tiers, positioning ERA-IT as a bridge between semantic IP analytics and transparent economic valuation while addressing both the opacity of black-box language models and the latency of traditional citation-based indicators.
@@ 2.3 Generative AI for IP Management
Recent work on instruction-tuned generative models provides the technical basis for structured reasoning in ERA-IT. Instruction tuning surveys indicate that LLMs can be further trained on collections of instruction-response pairs to improve controllability and domain adaptation while systematising common pitfalls of instruction-driven alignment zhang2023 . Chain-of-Thought prompting studies demonstrate that demonstrations with explicit intermediate steps can markedly improve multi-step reasoning, and that relevance and step ordering matter more than strictly correct logical derivations wang2023cot . Research on self-explanations finds that model-generated rationales may appear plausible yet diverge from traditional attribution methods, motivating explicit checks of explanation faithfulness huang2023selfexplain . In high-stakes legal settings, participatory design with practising lawyers argues that LLMs should primarily function as decision-support tools within explicit policy frameworks rather than as autonomous legal advisers cheong2024law . Within the IP domain, domain-specific frameworks guide models along patent-relevant dimensions. For example, PatentMind uses a multi-aspect reasoning graph to decompose and score patent similarity across technical, legal, and contextual dimensions yoo2025patentmind ; PatentScore defines a multi-dimensional rubric for evaluating LLM-generated claims across structural, semantic, and legal aspects yoo2025patentscore ; and Self Filtered Distillation uses LLM-generated labels and rationales to compute trust scores, thereby filtering or reweighting training examples for patent classification yoo2025sfd .
@@
However, from the perspective of ex-ante patent valuation, these techniques remain only loosely connected to economic ground truth. Instruction tuning and Chain-of-Thought prompting are typically optimised on general reasoning benchmarks and do not directly constrain rationales or predictions with observable economic outcomes like renewal behaviour, so their explanations are at best indirectly related to value signals zhang2023 ; wang2023cot . Studies of self-explanations further show that natural language rationales can appear faithful under coarse metrics while assigning very different importance patterns, which complicates their use as evidence for economic reasoning huang2023selfexplain . In the patent domain, PatentMind and PatentScore align model behaviour with expert assessments of similarity and claim quality, and Self Filtered Distillation improves the reliability of patent classifiers using trust indicators derived from model outputs, yet none of these frameworks explicitly tie reasoning to renewal trajectories or other revealed preference signals yoo2025patentmind ; yoo2025patentscore ; yoo2025sfd . Therefore, ERA-IT couples patent renewal history, treated as revealed economic preference, with a four-step Economic Chain-of-Thought schema covering technical strength, market context, legal robustness, and value synthesis, and uses teacher-guided data synthesis together with parameter-efficient adaptation to align LLM reasoning with these economic signals and to turn generated explanations into actionable features for ex-ante patent valuation.
3 Dataset Construction

### v2

@@ Abstract
This study proposes the Economic Reasoning Alignment via Instruction Tuning (ERA-IT) framework, which aligns the semantic reasoning of Large Language Models (LLMs) with revealed economic preferences to establish a new paradigm for real-time and interpretable patent valuation. We theoretically conceptualize patent renewal history as a revealed economic preference and leverage it as an objective supervisory signal to align the generative reasoning of LLMs with market realities, a process we term Eco-Semantic Alignment. Our empirical analysis of 10,000 randomly sampled European Patent Office patents across diverse technological domains confirms the efficacy of this approach. The ERA-IT framework achieves superior performance across all evaluated metrics compared to the baseline models. Specifically, it surpasses the traditional Random Forest model based on TF-IDF features by 36.3% in accuracy, 33.1% in Macro-F1, and 88.1% in the Matthews Correlation Coefficient (MCC). Against the discriminative pre-trained language model Longformer, ERA-IT exhibits gains of 11.5% in accuracy, 8.7% in Macro-F1, and 27.4% in MCC. It also exceeds the zero-shot performance of the generative LLM GPT-5-mini by 9.6% in accuracy, 8.3% in Macro-F1, and 21.5% in MCC. The ablation study confirms the criticality of the Rationale Generation module (Economic Chain-of-Thought), whose removal led to the largest performance decline, with a 6.4% drop in Macro-F1. Moreover, ERA-IT demonstrates robust applicability across heterogeneous patent landscapes.
@@ 1 Introduction
In the contemporary knowledge-driven economy, patents, as core intangible assets, are increasingly pivotal to firms’ strategic decisions and value creation (ernst2003). They serve as the fundamental currency of innovation, delineating a firm’s technological trajectory and securing market dominance (hall2005; reitzig2004). Accurate and timely patent valuation forms the bedrock of firms’ strategy for managing technological innovation (fischer2014; kalip2022). However, achieving objective and accurate ex-ante valuation of patents remains a classic and persistent grand challenge in information science, information economics, and technological innovation management (reitzig2004). Figure 1 outlines the evolutionary path of patent valuation methodologies.
@@
The crux of this challenge stems from two critical dimensions. First, patents themselves are complex and information-dense hybrid “technical-legal” artifacts. Their true value is deeply encoded in unstructured technical texts (han2015). This creates profound information asymmetry between informed patent holders and external evaluators, including investors and managers, all of whom struggle to penetrate this “technical black box” and transact efficiently in markets for technology (akerlof1970; gambardella2007; gans2008). This opacity creates severe market friction, where stakeholders struggle to distinguish high-value inventions from the vast noise of global filings due to the high-dimensional complexity of mapping technical specifications to market outcomes (pakes1986; trajtenberg1990). Second, information asymmetry is inherent in the global patent ecosystem. Unlike domain-specific studies that benefit from homogeneous terminologies, a generalized valuation framework must contend with heterogeneous linguistic patterns across disparate technological sectors (pakes1984). In this context, traditional bibliometric indicators, such as citation counts, fail to serve as effective signals for timely decision-making due to systemic temporal latency; a patent typically accumulates meaningful citation data only years after its publication (harhoff2003). This “backward-looking” nature of bibliometrics prevents managers from assessing patent value in a way that is both real-time and closely reflects genuine market preferences. Consequently, there exists an urgent imperative to bridge the gap between technical disclosure and economic realization. This imperative drives the quest for advanced architectures capable of leveraging “forward-looking” signals, constituting a critical research frontier.
@@
In this context, historical data on patent renewal has gained theoretical prominence as a direct signal of “revealed economic preferences,” since the act of renewal represents a “vote with the wallet” that is more closely aligned with market reality than any external proxy (pakes1986; pakes1989; schankerman1986). From the perspective of economic signaling theory (jiang2026), the decision to pay monotonically increasing renewal fees reveals the patent holder’s preference, indicating that the asset’s private value exceeds its maintenance expenses (danish2020; jou2018; porter2023; kumar2024). Unlike stochastic citations, renewal records provide a definitive, discrete signal of value maintenance available throughout the patent’s lifecycle. However, a fundamental methodological gap persists: how to effectively align this structured, low-dimensional economic-behavior signal with the high-dimensional, unstructured semantic information in patent texts to build an accurate and explainable predictive model. Existing frameworks have struggled to integrate this objective economic signal with the rich, unstructured semantic information embedded within patent documents (arts2018). Traditional Natural Language Processing (NLP) techniques often fail to capture the nuanced causal relationships between the specific exclusionary scope defined in the claims and the patent’s potential for economic longevity (lee2020).
@@
The recent advent of Large Language Models (LLMs) offers a transformative opportunity to overcome these linguistic barriers (yan2025). LLMs possess the unique capacity for complex reasoning and for extracting latent semantic features. Nevertheless, applying general-purpose LLMs to patent valuation is non-trivial; these models often lack the domain-specific intuition required to align technical novelty with economic viability. To overcome this critical barrier in high-stakes Intellectual Property (IP) management, there is an urgent need for adaptive technological and methodological innovations. The development of a predictive architecture that effectively fuses the objective ground truth of renewal data with the semantic reasoning capabilities of LLMs, a process that we term “Eco-Semantic Alignment,” represents a largely unexplored frontier in intelligent information systems.
@@
To bridge this methodological gap, this study proposes the Economic-Reasoning Alignment via Instruction Tuning (ERA-IT) framework, as illustrated in Figure 2. Our approach uses European Patent Office (EPO) renewal data as a supervisory signal to guide LLMs’ generative reasoning. Guided by a rigorous data fusion strategy, we operationalize the longitudinal payment of renewal fees as a value proxy. This allows us to translate economic behavior into a labeled dataset with discrete value tiers, transforming the ambiguous task of valuation into a robust supervised learning problem. Subsequently, we employ a strategy that synergizes parameter-efficient fine-tuning with chain-of-thought prompting. This alignment enables the model to learn latent semantic patterns correlated with economic survival, effectively converting unstructured text into quantifiable predictive indicators.
To rigorously validate the efficacy and generalizability of this framework, we conducted a comprehensive empirical analysis using a dataset comprising over 10,000 EPO patents, randomly sampled across various International Patent Classification (IPC) sections. This randomized sampling strategy ensures that our findings are not biased toward specific high-velocity domains and instead represent a robust solution applicable to the broader technological landscape. Our results demonstrate that this hybrid approach yields significantly superior predictive accuracy compared to conventional rule-based heuristics and statistical baselines (arts2018). By minimizing predictive latency in citation metrics, our system provides a scalable solution for automated high-precision IP valuation.
Our study pioneers a theoretical framework that aligns LLMs’ semantic reasoning with revealed economic preferences to establish a new paradigm for real-time and interpretable patent valuation. It has at least threefold research contributions, summarized as follows:
@@
  • Methodological Contribution in Reasoning Alignment: We establish the algorithmic efficacy of the ERA-IT architecture, which synergizes parameter-efficient fine-tuning with domain-specific instruction tuning strategies. We provide empirical evidence that standard pre-trained models are insufficient for the specialized task of IP valuation, demonstrating that our domain-adaptation process aligns the model’s semantic processing with the “Economic Chain-of-Thought” underlying obfuscated legal-technical texts.
@@ 2.1 Patent Valuation
Research on patent valuation has shifted from simple patent counts toward richer indicators that proxy the economic and innovative impact of inventions (kalip2025; schmitt2025). Using the American Inventors Protection Act, which mandated earlier publication of most US patent applications, liu2023 explore a multi-task learning based identification model to identify the high-value patents and the standard-essential patents. hegde2023 show that earlier disclosure is associated with more and faster follow-on citations, fewer abandonments, and higher R&D investment, highlighting how publication rules shape both the private and social value of patents. svensson2022 analyzes four widely used patent value indicators, renewal length, forward citations, family size, and documents that each capture distinct facets of technological and economic value rather than a single underlying measure. Building on this indicator-based tradition, ponta2021 propose the Innovation Patent Index (IPI), a composite innovation-performance measure that aggregates multiple patent characteristics related to technological quality, market potential, and scope into a firm-level index. Complementing these firm- and patent-level perspectives, park2023 apply the previously introduced CD index of disruption to large-scale datasets of papers and patents, extend it with normalized variants and alternative measures, and show that science and technology have become less disruptive over time, illustrating the potential of advanced citation-based metrics for characterizing the qualitative role of inventions in the knowledge network. lee2025 proposes a machine learning framework based on structured bibliometrics and inventor-related indicators. By introducing three underexplored value metrics and applying six algorithms across five distinct strategies to predict the value of renewable energy patents, the findings demonstrate that random forest and artificial neural network models, which integrate multiple indicator categories, achieve the highest accuracy.
However, from the perspective of ex-ante patent valuation, these approaches remain largely ex-post and rely on outcome variables that materialize only after a substantial time lag (fu2025; liu2024a). Renewal behavior, composite indices such as IPI, and disruption-based citation metrics are typically calibrated on long-term renewal patterns, realized commercialization, or accumulated citation structures, and they make limited use of the full semantic content of patent texts at the level of individual inventions (svensson2022; ponta2021; park2023). Moreover, these indicators generally summarize value at the firm or field level and do not provide explicit, stepwise economic rationales for why a particular patent should be assessed as more or less valuable at the time of filing. Because of these limitations, existing patent-valuation approaches remain constrained in their ability to provide economically grounded and explainable ex-ante value assessments for newly filed patents. These limitations underscore the need for ex-ante patent valuation frameworks that integrate semantic text analysis with interpretable reasoning, such as ERA-IT.
@@ 2.2 Semantic Analysis for IP Management
Recent work in IP management increasingly relies on semantic analysis of patent and related technical text (jiang2025). ali2024 review patent prior-art retrieval, tracing the field’s progression from basic Boolean and keyword search and classic information retrieval models toward semantic retrieval enabled by NLP and deep learning. xu2025 extend this line of work to industry–academia collaboration by proposing a semantic retrieval approach that uses text embeddings to analyze and rank collaborative research outputs. yun2022 propose a method for mining technology trends based on patent semantic analysis. Through a four-stage process involving the extraction of technical purposes and effects, clustering-based lexicon construction, mining of evolutionary patterns, and forecasting future directions by referencing models from other fields, it supports enterprises in identifying new technology opportunities. At the model level, chikkamath2022 construct large-scale USPTO subclass datasets and investigate how a BERT for Patents encoder and strong non-neural baselines use claims and abstracts for multi-label patent classification. Complementary research on structured representations decomposes Chinese patent claims into function, structure, and purpose segments and applies pre-trained transformer-based models to classify them into technical categories, thereby improving classification accuracy and supporting retrieval of patents in similar technical domains (li2025). In parallel, zini2022 survey explainability techniques for deep NLP models, cataloguing model-agnostic and model-specific methods that interpret embeddings, internal representations, and output decisions for text-based tasks. yoshikawa2025 enhance automatic patent classification by leveraging GPT-3.5-turbo to generate summaries from patent claims and detailed descriptions, which are then used to fine-tune RoBERTa (for English) and XLM-RoBERTa (for Japanese) models.
@@
Despite these advances, existing semantic approaches are optimised for retrieval, classification, or document alignment rather than for transparent ex-ante economic valuation of patents. The patent retrieval survey focuses on relevance and recall, leaving the relationship between semantic similarity and economic value largely implicit (ali2024). Semantic matching systems for industry–academia collaboration operate in high-dimensional embedding spaces but do not attempt to infer monetary or strategic value or to incorporate ground-truth financial signals (xu2025). Models such as the BERT for Patents and the structured claim classifiers in the Chinese patent study are trained to predict technology codes, not renewal-based value tiers, and typically output numerical labels without explicit economic rationales (chikkamath2022; li2025; abbasiantaeb2025). Even where explainability is the primary topic, current surveys emphasise post-hoc interpretation tools for generic NLP tasks and note that rigorous evaluation in high-stakes legal and economic contexts remains limited (zini2022). In contrast, ERA-IT explicitly fuses semantic reasoning with revealed preference renewal data and uses instruction-tuned chain-of-thought generation to map claim language to value tiers, positioning ERA-IT as a bridge between semantic IP analytics and transparent economic valuation while addressing both the opacity of black-box language models and the latency of traditional citation-based indicators.
@@ 2.3 Generative AI for IP Management
Recent work on instruction-tuned generative models provides the technical basis for structured reasoning in ERA-IT. Instruction tuning surveys indicate that LLMs can be further trained on collections of instruction-response pairs to improve controllability and domain adaptation while systematising common pitfalls of instruction-driven alignment (zhang2023). Chain-of-Thought prompting studies demonstrate that demonstrations with explicit intermediate steps can markedly improve multi-step reasoning, and that relevance and step ordering matter more than strictly correct logical derivations (wang2023cot). Research on self-explanations finds that model-generated rationales may appear plausible yet diverge from traditional attribution methods, motivating explicit checks of explanation faithfulness (huang2023selfexplain). In high-stakes legal settings, participatory design with practising lawyers argues that LLMs should primarily function as decision-support tools within explicit policy frameworks rather than as autonomous legal advisers (cheong2024law). xiong2025 propose a large language model-assisted active learning framework designed for scalable multi-label patent classification, validated on a dataset of 100,000 drone-related patents. By integrating GPT-4’s uncertainty estimation with diversity-aware sampling, this approach significantly reduces the need for manual annotation while maintaining strong classification performance and improving key evaluation metrics. Within the IP domain, domain-specific frameworks guide models along patent-relevant dimensions (shomee2025). For example, PatentMind uses a multi-aspect reasoning graph to decompose and score patent similarity across technical, legal, and contextual dimensions (yoo2025patentmind); PatentScore defines a multi-dimensional rubric for evaluating LLM-generated claims across structural, semantic, and legal aspects (yoo2025patentscore); and Self Filtered Distillation uses LLM-generated labels and rationales to compute trust scores, thereby filtering or reweighting training examples for patent classification (yoo2025sfd).
@@
However, from the perspective of ex-ante patent valuation, these techniques remain only loosely connected to economic ground truth. Instruction tuning and Chain-of-Thought prompting are typically optimised on general reasoning benchmarks and do not directly constrain rationales or predictions with observable economic outcomes like renewal behaviour, so their explanations are at best indirectly related to value signals (zhang2023; wang2023cot). Studies of self-explanations further show that natural language rationales can appear faithful under coarse metrics while assigning very different importance patterns, which complicates their use as evidence for economic reasoning (huang2023selfexplain). In the patent domain, PatentMind and PatentScore align model behaviour with expert assessments of similarity and claim quality, and Self Filtered Distillation improves the reliability of patent classifiers using trust indicators derived from model outputs, yet none of these frameworks explicitly tie reasoning to renewal trajectories or other revealed preference signals (yoo2025patentmind; yoo2025patentscore; yoo2025sfd). Therefore, ERA-IT couples patent renewal history, treated as revealed economic preference, with a four-step Economic Chain-of-Thought schema covering technical strength, market context, legal robustness, and value synthesis, and uses teacher-guided data synthesis together with parameter-efficient adaptation to align LLM reasoning with these economic signals and to turn generated explanations into actionable features for ex-ante patent valuation.
3 Research Objectives
We explore three critical research questions underexplored in prior studies:
  • RQ1 (Efficacy): Does ERA-IT outperform state-of-the-art discriminative baselines and LLMs in predicting patent value?
  • RQ3 (Generalizability): Is the framework robust across heterogeneous technological domains (IPC sections) and low-resource settings?
Based on the proposed framework and the defined research inquiries, this study pursues three specific objectives: (1) To empirically substantiate the predictive efficacy of the ERA-IT framework against state-of-the-art discriminative baselines and zero-shot generative LLMs for patent value prediction, using patent renewal history as the ground-truth signal. This objective directly tests the core claim that aligning LLM reasoning with revealed economic preferences, a process known as Eco-Semantic Alignment, yields superior and timely performance metrics. (2) To deconstruct and explain the contributory mechanism of explicit economic rationale generation. Through ablation studies and comparative analysis, this objective aims to isolate and quantify the impact of the Economic Chain-of-Thought module, thereby elucidating how logically grounded rationales enhance model fidelity and serve as a transparent cognitive scaffold for decision-making. (3) To rigorously verify the robustness and generalizability of the framework across heterogeneous technological domains classified by IPC sections and under low-resource conditions. This objective evaluates the practical utility and scalability of ERA-IT, ensuring its reliability beyond the training distribution and confirming its applicability to diverse innovation management contexts.
4 Dataset